The equation of a line passing through two distinct points (x1,y1) and (x2,y2) is given by
x y 1
x2 y2 1 =0
x3 y3 1
Use the determinant to write an equation for the line passing through (-4,-4) and (-8,9). Express the lines equation in standard form.
a. 9x-4y+32=0
b. -4x-8y-36=0
c. 13x+4y-68=0
d.-13x-4y-68=0

Answer by Alan3354(69443)

(-4,-4) and (-8,9).
------------------
|x y +1|
|-4-4 1| = 0
|-8 9 1|
-------------
x(-4-9) - y(-4+8) + -36-32 = 0
-13x - 4y = 68 Answer d
13x + 4y = -68